# Operator World Models for Reinforcement Learning

### v1

@@ 3 Proposed Estimator and Algorithm
Under Asm. 1, $({T}|_{\mathcal{F}})^{*}=({T}^{*})|_{\mathcal{G}}$ and, for any $(x,a)\in\Omega$
$({T}|_{\mathcal{F}})^{*}\psi(x,a)=\int_{\mathcal{X}}\varphi(x^{\prime})\leavevmode\nobreak\ \tau(x^{\prime}|x,a)=E[\varphi(X^{\prime})|X=x,A=a].$ (9)
Prop. 2 (see proof in Sec. A.2) shows that 9 is equivalent to the standard linear MDP assumption (see e.g. [26, Ch. 8]) when $\mathcal{X}$ is a finite set (taking $\varphi$ the one-hot encoding) while being weaker in infinite settings. From the CME perspective, the proposition characterizes the action of $({T}|_{\mathcal{F}})^{*}$ as sending evaluation vectors in ${\mathcal{G}}$ to the conditional expectation of evaluation vectors in ${\mathcal{F}}$ with respect to $\tau$, the definition of conditional mean embedding of $\tau$ [27, 15]. This characterization also suggests a learning strategy: 9 characterizes the action of ${T}$ as evaluating to the conditional expectation of a vector $\varphi(X^{\prime})$ given $(x,a)$. Given a set of points $(x_{i},a_{i})_{i=1}^{n}$ and corresponding $x^{\prime}$ sampled from $\tau(\cdot|x_{i},a_{i})$, this can be learned by minimizing the squared loss, yielding the estimator
${T}_{n}=S_{n}^{*}(K_{n}+n\lambda{Id})^{-1}Z_{n}=\operatorname*{argmin}_{{T}\in{HS}({\mathcal{F}},{\mathcal{G}})}\leavevmode\nobreak\ \frac{1}{n}\sum_{i=1}^{n}\left\|{\varphi(x_{i}^{\prime})-{T}^{*}\psi(x_{i},a_{i})}\right\|_{\mathcal{F}}^{2}+\lambda\left\|{{T}}\right\|_{{HS}}^{2}.$ (10)
When ${\mathcal{F}}$ and ${\mathcal{G}}$ are finite dimensional, $S_{n}$ and $Z_{n}$ are matrices with $n$ rows, each corresponding respectively to the vectors $\psi(x_{i},a_{i})$ and $\varphi(x_{i}^{\prime})$ for $i=1,\dots,n$. In the infinite setting they generalize to operators $S_{n}:{\mathcal{G}}\to\mathbb{R}^{n}$ and $Z_{n}:{\mathcal{F}}\to\mathbb{R}^{n}$. The $K_{n}=S_{n}S_{n}^{*}\in\mathbb{R}^{n\times n}$ is the Gram (or kernel) matrix with $(i,j)$-th entry corresponding to $(K_{n})_{ij}=\left\langle{\psi(x_{i},a_{i})},{\psi(x_{j},a_{j})}\right\rangle$. We conclude our discussion on learning world models via CMEs by noting that in most RL settings, the reward function is unknown, too. Analogously to what we have described for ${T}_{n}$ and following the standard practice in supervised settings, we can learn an estimator for $r$ solving a problem akin to 10. This yields a function of the form $r_{n}=S_{n}^{*}b=\sum_{i=1}^{n}b_{i}\,\psi(x_{i},a_{i})$ as the linear combination of the embedded training points with the entries of the vector $b=(K_{n}+n\lambda{Id})^{-1}y$ where $y\in\mathbb{R}^{n}$ is the vector with entries $y_{i}=r(x_{i},a_{i})$.
@@
We now propose our strategy to generate an estimator for the action-value function $q_{\pi}$ of a given policy $\pi$ in terms of an estimator for the reward $r$ and a world model for ${T}$ learned in terms of the restriction to ${\mathcal{G}}$ and ${\mathcal{F}}$. To this end, we need to introduce the notion of compatibility between a policy $\pi$ and the pair $({\mathcal{G}},{\mathcal{F}})$.
@@
A policy $\pi:\mathcal{X}\to{\mathcal{M}}_{1}^{+}(\mathcal{A})$ is compatible with two subspaces ${\mathcal{F}}\subseteq B_{b}(\mathcal{X})$ and ${\mathcal{G}}\subseteq B_{b}(\Omega)$ if the restriction ${P}_{\pi}$ to ${\mathcal{G}}$ is a bounded linear operator with range $\subseteq{\mathcal{F}}$, that is $({P}_{\pi})|_{\mathcal{G}}:{\mathcal{G}}\to{\mathcal{F}}$.
Definition 1 is analogous to the linear MDP Asm. 1 in that it requires the restriction of ${P}_{\pi}$ to ${\mathcal{G}}$ to take values in the associated space ${\mathcal{F}}$. However, it is slightly weaker since it requires this restriction to be bounded (and linear) rather than being a HS operator. We will discuss in Sec. 4 how this difference will allow us to show that a wide range of policies (in particular those generated by our POWR  method) is $({\mathcal{G}},{\mathcal{F}})$-compatible for our choice of function spaces. Definition 1 is the key condition that enables us to generate an estimator for $q_{\pi}$, as characterized by the following result.
@@
Let ${T}_{n}=S_{n}^{*}BZ_{n}\in{HS}({\mathcal{F}},{\mathcal{G}})$ and $r_{n}=S_{n}^{*}b\in{\mathcal{G}}$ for respectively a $B\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$. Let $\pi$ be $({\mathcal{G}},{\mathcal{F}})$-compatible. Then,
$\hat{q}_{\pi}=({Id}-\gamma{T}_{n}{P}_{\pi})^{-1}r_{n}=S_{n}^{*}({Id}-\gamma BM_{\pi n})^{-1}b$ (11)
where $M_{\pi n}=Z_{n}{P}_{\pi}S_{n}^{*}\in\mathbb{R}^{n\times n}$ is the matrix with entries
$(M_{\pi n})_{ij}=\int_{\mathcal{A}}\left\langle{\psi(x_{i}^{\prime},a)},{\psi(x_{j},a_{j})}\right\rangle\leavevmode\nobreak\ \pi(da|x_{i}^{\prime}).$ (12)
Prop. 3 leverages a kernel trick argument to express the estimator for the action-value function of $\pi$ as the linear combination $\hat{q}_{\pi}=\sum_{i=1}^{n}c_{i}\,\psi(x_{i},a_{i})$ of the (embedded) training points $\psi(x_{i},a_{i})$ and the entries $c_{i}$ of the vector $c=(I-\gamma BZ_{n}{P}_{\pi}S_{n}^{*})^{-1}b\in\mathbb{R}^{n}$. We prove the result in Sec. A.4. We note that in 11 both $B$ and $Z_{n}{P}_{\pi}S_{n}^{*}$ are $n\times n$ matrices and therefore the characterization of $\hat{q}_{\pi}$ amounts to solving a $n\times n$ linear system. For settings where $n$ is large, one can adopt random projection methods such as Nyström approximation to learn ${T}_{n}$ and $r_{n}$ [28]. These strategies have been recently shown to significantly reduce the computational load of learning while retaining the same empirical and theoretical performance as their non-approximated counterparts [29, 30].
We conclude this section noting how 12 implies that we only need to be able to evaluate $\pi$, but we do not need explicit knowledge of ${P}_{\pi}$ as operator. As we shall see, this property will be instrumental to prove generalization bounds for our proposed PMD algorithm in Sec. 4.
@@
  Input: Dataset $\mathcal{D}=(x_{i},a_{i},x_{i}^{\prime},r_{i})_{i=1}^{n}$, discount $\gamma\in(0,1)$, step size $\eta>0$, kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ initial weights $C_{0}=0\in\mathbb{R}^{n\times|\mathcal{A}|}$
  Let $E\in\mathbb{R}^{n\times|\mathcal{A}|}$ with rows $E_{i}=$ $\text{\sc OneHot}_{|\mathcal{A}|}(a_{i})$ for $i=1,\dots,n$.
  Let $K,H\in\mathbb{R}^{n\times n}$ such that $K_{ij}=k(x_{i},x_{j})\delta_{a_{i}=a_{j}}$ and $H_{ij}=k(x_{i}^{\prime},x_{j})$ $\forall i,j=1,\dots,n$.
  Let $B=(K+\lambda{Id})^{-1}$ and $b=By$ with $y=(r_{1},\dots,r_{n})\in\mathbb{R}^{n}$
  For $t=0,1,\dots,T$:
    $\pi=$ softmax$(\eta Q)$ with $Q=HC_{t}$
    Let $M\in\mathbb{R}^{n\times n}$ such that $M_{ij}=H_{ij}\pi_{ja_{i}}$ for $i,j=1,\dots,n$
    $C_{t+1}=C_{t}+$diag$(c)E$ with $c=({Id}-\gamma BM)^{-1}b$
  Return $\pi_{T}:\mathcal{X}\to\Delta(\mathcal{A})$ such that $\pi_{T}(x)=$softmax$(\eta\,f(x))$ with $f(x)=k(x)C_{T}$
  Return and $k(x)=(k(x,x_{1}),\dots,k(x,x_{n}))$ for all $x\in\mathcal{X}$
@@ 4 Proposed Algorithm: POWR
In POWR, we combine the CME world model presented in Sec. 3 with mirror descent steps using the Kullback-Leibler divergence $D_{{\rm KL}}(p;p^{\prime})=\sum_{a\in\mathcal{A}}p_{a}\log(p_{a}/p_{a}^{\prime})$ in the update of 2. It was shown in [10] that in this case PMD corresponds to Natural Policy Gradient [24], while [31, Example 9.10] shows how the solution to 2 can be written in closed form for any $x\in\mathcal{X}$ as
$\pi_{t+1}(\cdot|x)=\frac{\pi_{t}(\cdot|x)e^{\eta\hat{q}_{\pi_{t}}(x,\cdot)}}{\sum_{a\in\mathcal{A}}\pi_{t}(a|x)e^{\eta\hat{q}_{\pi_{t}}(x,a)}}.$ (13)
Additionally, the formula above can be applied recursively, expressing $\pi_{t+1}$ as the softmax operator applied to the discounted sum of the action-value functions up to the current iteration
@@
A key question to address to adopt the action-value estimators introduced in Sec. 3 is choosing the two spaces ${\mathcal{F}}$ and ${\mathcal{G}}$ to perform world model learning. Specifically, to apply Prop. 3 and obtain proper estimators $\hat{q}_{\pi_{t}}$, we need to guarantee that all policies generated by the PMD update are $({\mathcal{G}},{\mathcal{F}})$-compatible (Definition 1). The following result describes a suitable family of such spaces.
@@
Prop. 4 (proof in Sec. A.5) states that for the specific choice of function spaces ${\mathcal{F}}={\mathcal{H}}\otimes{\mathcal{H}}$ and ${\mathcal{G}}=\mathbb{R}^{|\mathcal{A}|}\otimes{\mathcal{H}}$, we can guarantee $({\mathcal{G}},{\mathcal{F}})$-compatibility, provided that ${\mathcal{H}}$ is rich enough to “contain” all $\pi(a|\cdot)$ for $a\in\mathcal{A}$. We postpone the discussion on identifying a suitable spaces ${\mathcal{H}}$ for PMD to Sec. 5, since $({\mathcal{G}},{\mathcal{F}})$-compatibility is not needed to mechanically apply Prop. 3 and obtain an estimator $\hat{q}_{\pi}$. This is because 11 exploits a kernel-trick to bypass the need to know ${P}_{\pi}$ explicitly and rather requires only to be able to evaluate $\pi$ on the training data. The latter is possible for $\pi_{t+1}$, thanks to the point-wise characterization of the PMD update in 13. We can therefore present our algorithm.
@@
Alg. 1 describes Policy mirror descent with Operator World-models for Reinforcement learning (POWR). Following the intuition of Prop. 4, the algorithm assumes to work with separable spaces. During an initial phase, we learn the world model ${T}_{n}=S_{n}^{*}BZ_{n}$ and the reward $r_{n}=S_{n}^{*}b$ fitting the conditional mean embedding described in 10 on a dataset $(x_{i},a_{i},x_{i}^{\prime},r_{i})_{i=1}^{n}$ (e.g. obtained via experience replay [32]). Once the world model has been learned, we optimize the policy and perform PMD iterations via 14. In this second phase, we first evaluate the past (cumulative) action-value estimators $\sum_{s=0}^{t}\hat{q}_{s}$ to obtain the current policy $\pi_{t+1}(\cdot|x_{i})$ by (inexact) PMD via the softmax operator in 14. We use the newly obtained policy to compute the matrix $M_{\pi_{t+1}n}$ defined in 12, which is a key component to obtain the estimator $\hat{q}_{t+1}$ for $q_{\pi_{t+1}}$. We note that in the case of the separable spaces of Prop. 4, this matrix reduces to the $n\times n$ matrix $M\in\mathbb{R}^{n\times n}$ with entries $M_{ij}=k(x_{i}^{\prime},x_{j})\pi_{t+1}(a_{j}|x_{i}^{\prime})$, where $k(x_{i}^{\prime},x_{j})=\left\langle{\phi(x_{i}^{\prime})},{\phi(x_{j})}\right\rangle$. Finally, we obtain $c=({Id}-\gamma BM)^{-1}b$ and model $\hat{q}_{t+1}=S_{n}^{*}c$ according to Prop. 3.
Clearly, world model learning phase and PMD can be alternated in POWR , essentially finding a trade-off between exploration and exploitation. This could possibly lead to a refinement of the world model as more observations are integrated into the estimator. While in this work we do not investigate the effects of such alternating strategy, Thm. 7 offers relevant insights in this sense. The result characterizes the behavior of the PMD algorithm when combined with a varying (possibly increasing) accuracy in the estimation of the action-valued function (see Sec. 5 for more details).

### v2

@@ 3 Operator World Models
Under Asm. 1, $(\mathsf{T}|_{\mathcal{F}})^{*}=(\mathsf{T}^{*})|_{\mathcal{G}}$ and, for any $(x,a)\in\Omega$
$(\mathsf{T}|_{\mathcal{F}})^{*}\psi(x,a)=\int_{\mathcal{X}}\varphi(x^{\prime})\leavevmode\nobreak\ \tau(x^{\prime}|x,a)=\mathbb{E}[\varphi(X^{\prime})|X=x,A=a].$ (9)
Proposition 2 shows that 9 is equivalent to the standard linear MDP assumption [30, Ch. 8] when $\mathcal{X}$ is a finite set (taking $\varphi$ the one-hot encoding) while being weaker in infinite settings. From the CME perspective, the proposition characterizes the action of $(\mathsf{T}|_{\mathcal{F}})^{*}$ as sending evaluation vectors in ${\mathcal{G}}$ to the conditional expectation of evaluation vectors in ${\mathcal{F}}$ with respect to $\tau$, the definition of conditional mean embedding of $\tau$ [31, 15]. This characterization also suggests a learning strategy: 9 characterizes the action of $\mathsf{T}$ as evaluating the conditional expectation of a vector $\varphi(X^{\prime})$ given $(x,a)$. Given a set of points $(x_{i},a_{i})_{i=1}^{n}$ and corresponding $x^{\prime}$ sampled from $\tau(\cdot|x_{i},a_{i})$, this can be learned by minimizing the squared loss, yielding the estimator (see  [15, Sec 4.2])
$\mathsf{T}_{n}=\operatornamewithlimits{argmin}_{\mathsf{T}\in\mathsf{HS}({\mathcal{F}},{\mathcal{G}})}\leavevmode\nobreak\ \frac{1}{n}\sum_{i=1}^{n}\left\|{\varphi(x_{i}^{\prime})-\mathsf{T}^{*}\psi(x_{i},a_{i})}\right\|_{\mathcal{F}}^{2}+\lambda\left\|{\mathsf{T}}\right\|_{\mathsf{HS}}^{2}=S_{n}^{*}K_{\lambda}^{-1}Z_{n}.$ (10)
When ${\mathcal{F}}$ and ${\mathcal{G}}$ are finite dimensional, $S_{n}$ and $Z_{n}$ are matrices with $n$ rows, each corresponding respectively to the vectors $\psi(x_{i},a_{i})$ and $\varphi(x_{i}^{\prime})$ for $i=1,\dots,n$. In the infinite setting, they generalize to operators $S_{n}:{\mathcal{G}}\to{\mathbb{R}}^{n}$ and $Z_{n}:{\mathcal{F}}\to{\mathbb{R}}^{n}$. The matrix $K_{\lambda}=S_{n}S_{n}^{*}+n\lambda\mathsf{Id}_{n}\in{\mathbb{R}}^{n\times n}$ is the regularized Gram (or kernel) matrix with $(i,j)$-th entry corresponding to
$\big{(}K_{\lambda}\big{)}_{ij}=\left\langle{\psi(x_{i},a_{i})},{\psi(x_{j},a_{j})}\right\rangle+n\lambda\delta_{ij}.$ (11)
We conclude our discussion on learning world models via CMEs by noting that in most RL settings, the reward function is unknown, too. Analogously to what we have described for $\mathsf{T}_{n}$ and following the standard practice in supervised settings, we can learn an estimator for $r$ solving a problem akin to 10. This yields a function of the form $r_{n}=S_{n}^{*}b=\sum_{i=1}^{n}b_{i}\,\psi(x_{i},a_{i})$ as the linear combination of the embedded training points with the entries of the vector $b=K_{\lambda}^{-1}y$ where $y\in{\mathbb{R}}^{n}$ is the vector with entries $y_{i}=r(x_{i},a_{i})$.
@@
We now propose our strategy to generate an estimator for the action-value function $q_{\pi}$ of a given policy $\pi$ in terms of an estimator for the reward $r$ and a world model for $\mathsf{T}$ learned in terms of the restriction to ${\mathcal{G}}$ and ${\mathcal{F}}$. To this end, we need to introduce the notion of compatibility between a policy $\pi$ and the pair $({\mathcal{G}},{\mathcal{F}})$.
@@
A policy $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{A})$ is compatible with two subspaces ${\mathcal{F}}\subseteq B_{b}(\mathcal{X})$ and ${\mathcal{G}}\subseteq B_{b}(\Omega)$ if the restriction $\mathsf{P}_{\pi}$ to ${\mathcal{G}}$ is a bounded linear operator with range $\subseteq{\mathcal{F}}$, that is $(\mathsf{P}_{\pi})|_{\mathcal{G}}:{\mathcal{G}}\to{\mathcal{F}}$.
Definition 1 is analogous to the linear MDP Asm. 1 in that it requires the restriction of $\mathsf{P}_{\pi}$ to ${\mathcal{G}}$ to take values in the associated space ${\mathcal{F}}$. However, it is slightly weaker since it requires this restriction to be bounded (and linear) rather than being an HS operator. We will discuss in Sec. 4 how this difference will allow us to show that a wide range of policies (in particular those generated by our POWR method) is $({\mathcal{G}},{\mathcal{F}})$-compatible for our choice of function spaces. Definition 1 is the key condition that enables us to generate an estimator for $q_{\pi}$, as characterized by the following result.
@@
Let $\mathsf{T}_{n}=S_{n}^{*}BZ_{n}\in\mathsf{HS}({\mathcal{F}},{\mathcal{G}})$ and $r_{n}=S_{n}^{*}b\in{\mathcal{G}}$ for respectively a $B\in{\mathbb{R}}^{n\times n}$ and $b\in{\mathbb{R}}^{n}$. Let $\pi$ be $({\mathcal{G}},{\mathcal{F}})$-compatible. Then,
$\hat{q}_{\pi}=(\mathsf{Id}-\gamma\mathsf{T}_{n}\mathsf{P}_{\pi})^{-1}r_{n}=S_{n}^{*}(\mathsf{Id}-\gamma BM_{\pi})^{-1}b$ (12)
where $M_{\pi}=Z_{n}\mathsf{P}_{\pi}S_{n}^{*}\in{\mathbb{R}}^{n\times n}$ is the matrix with entries
$\big{(}M_{\pi}\big{)}_{ij}=\left\langle{\varphi(x_{i}^{\prime})},{\mathsf{P}_{\pi}\psi(x_{j},a_{j})}\right\rangle=\int_{\mathcal{A}}\left\langle{\psi(x_{i}^{\prime},a)},{\psi(x_{j},a_{j})}\right\rangle\leavevmode\nobreak\ \pi(da|x_{i}^{\prime}).$ (13)
Proposition 3 leverages a kernel trick argument to express the estimator for the action-value function of $\pi$ as the linear combination $\hat{q}_{\pi}=\sum_{i=1}^{n}c_{i}\,\psi(x_{i},a_{i})$ of the (embedded) training points $\psi(x_{i},a_{i})$ and the entries $c_{i}$ of the vector $c=(\mathsf{Id}-\gamma BM_{\pi})^{-1}b\in{\mathbb{R}}^{n}$. We prove the result in Sec. A.4. We note that in 12 both $B$ and $M_{\pi}$ are $n\times n$ matrices and therefore the characterization of $\hat{q}_{\pi}$ amounts to solving a $n\times n$ linear system. For settings where $n$ is large, one can adopt random projection methods such as Nyström approximation to learn $\mathsf{T}_{n}$ and $r_{n}$ [32]. These strategies have been recently shown to significantly reduce the computational load of learning while retaining the same empirical and theoretical performance as their non-approximated counterparts [33, 34].
We conclude this section noting how 13 implies that we only need to be able to evaluate $\pi$, but we do not need explicit knowledge of $\mathsf{P}_{\pi}$ as operator. As we shall see, this property will be instrumental to prove generalization bounds for our proposed PMD algorithm in Sec. 4.
@@
Input: Dataset $(x_{i},a_{i},x_{i}^{\prime},r_{i})_{i=1}^{n}$, discount factor $\gamma\in(0,1)$, step size $\eta>0$, kernel function $k(x,x^{\prime})=\left\langle{\phi(x)},{\phi(x^{\prime})}\right\rangle$ with $\phi:\mathcal{X}\to{\mathcal{H}}$ as in Proposition 4, initial weights $C_{0}=0\in{\mathbb{R}}^{n\times|\mathcal{A}|}$.
/* World Model Learning */
let $E\in{\mathbb{R}}^{n\times|\mathcal{A}|}$ with rows $E_{i}=$ $\text{\sc OneHot}_{|\mathcal{A}|}(a_{i})$.
let $K_{\lambda}\in{\mathbb{R}}^{n\times n}$ such that $K_{ij}=k(x_{i},x_{j})\delta_{a_{i}=a_{j}}+n\lambda\delta_{ij}$triangleright Eq. 11
let $H\in{\mathbb{R}}^{n\times n}$ such that $H_{ij}=k(x_{i}^{\prime},x_{j})$
compute $K_{\lambda}^{-1}$ and $b=K_{\lambda}^{-1}y$ with $y=(r_{1},\dots,r_{n})\in{\mathbb{R}}^{n}$triangleright Eq. 10
/* Policy Mirror Descent */
for $t=0,1,\dots,T-1$ do:
       $\pi_{t+1}=$ softmax$(\eta HC_{t})\in{\mathbb{R}}^{n\times|\mathcal{A}|}$triangleright PMD Step (15)
       $M_{\pi_{t+1}}=H\odot(\pi_{t+1}E^{\top})\in{\mathbb{R}}^{n\times n}$triangleright Proposition 3, Eq. 13
       $C_{t+1}=C_{t}+$diag$(c)E$ with $c=(\mathsf{Id}-\gamma K_{\lambda}^{-1}M_{\pi_{t+1}})^{-1}b$triangleright Proposition 3, Eq. 12
end for
return $\pi_{T}:\mathcal{X}\to\Delta(\mathcal{A})$ such that $\pi_{T}(x)=$softmax$(\eta\,H_{x}C_{T})$ with $H_{x}=(k(x,x_{i}))_{i=1}^{n}\in{\mathbb{R}}^{n}$.
@@ 4 Proposed Algorithm: POWR
In POWR, we combine the CME world model presented in Sec. 3 with mirror descent steps using the Kullback-Leibler divergence $D_{{\rm KL}}(p;p^{\prime})=\sum_{a\in\mathcal{A}}p_{a}\log(p_{a}/p_{a}^{\prime})$ in the update of 2. It was shown in [10] that in this case PMD corresponds to Natural Policy Gradient [28]. As showed in [35, Example 9.10], the solution to 2 can be written in closed form for any $x\in\mathcal{X}$ as
$\pi_{t+1}(\cdot|x)=\frac{\pi_{t}(\cdot|x)e^{\eta\hat{q}_{\pi_{t}}(x,\cdot)}}{\sum_{a\in\mathcal{A}}\pi_{t}(a|x)e^{\eta\hat{q}_{\pi_{t}}(x,a)}},$ (14)
where we used the estimated action-value function $\hat{q}_{\pi}$ from Proposition 3. Additionally, the formula above can be applied recursively, expressing $\pi_{t+1}$ as the softmax operator applied to the discounted sum of the action-value functions up to the current iteration
@@
A key question to address before adopting the action-value estimators introduced in Sec. 3 is choosing the two spaces ${\mathcal{F}}$ and ${\mathcal{G}}$ to perform world model learning. Specifically, to apply Prop. 3 and obtain proper estimators $\hat{q}_{\pi_{t}}$, we need to guarantee that all policies generated by the PMD update (14) are $({\mathcal{G}},{\mathcal{F}})$-compatible (Def. 1). The following result describes a suitable family of such spaces.
@@
Proposition 4 (proof in Sec. A.5) states that for the specific choice of function spaces ${\mathcal{F}}={\mathcal{H}}\otimes{\mathcal{H}}$ and ${\mathcal{G}}={\mathbb{R}}^{|\mathcal{A}|}\otimes{\mathcal{H}}$, we can guarantee $({\mathcal{G}},{\mathcal{F}})$-compatibility, provided that ${\mathcal{H}}$ is rich enough to “contain” all $\pi(a|\cdot)$ for $a\in\mathcal{A}$. We postpone the discussion on identifying a suitable spaces ${\mathcal{H}}$ for PMD to Sec. 5, since $({\mathcal{G}},{\mathcal{F}})$-compatibility is not needed to mechanically apply Prop. 3 and obtain an estimator $\hat{q}_{\pi}$. This is because 12 exploits a kernel-trick to bypass the need to know $\mathsf{P}_{\pi}$ explicitly and rather requires only to be able to evaluate $\pi$ on the training data. The latter is possible for $\pi_{t+1}$, thanks to the explicit form of the PMD update in 14. We can therefore present our algorithm.
@@
[…]
Clearly, the world model learning and PMD phases can be alternated in POWR, essentially finding a trade-off between exploration and exploitation. This could possibly lead to a refinement of the world model as more observations are integrated into the estimator. While in this work we do not investigate the effects of such alternating strategy, Thm. 7 offers relevant insights in this sense. The result characterizes the behavior of the PMD algorithm when combined with a varying (possibly increasing) accuracy in the estimation of the action-valued function (see Sec. 5 for more details).